OpenStudy (anonymous):

nCn is __________ equal to 1. always sometimes never

OpenStudy (ash2326):

@Lunalovegood do you know how nCn is defined?

OpenStudy (anonymous):

How many different ways can you choose 3 marbles from a bag if there are 3 marbles in the bag?

OpenStudy (ash2326):

@SmoothMath that's a smooth answer:P

OpenStudy (anonymous):

Ahahaha thank ye kindly.

OpenStudy (anonymous):

no i don't @ash2326

OpenStudy (ash2326):

Do you know what is n! ?

OpenStudy (anonymous):

no..our teacher just kind of sprung this on us with out us knowing what it is

OpenStudy (ash2326):

Ok, n! is defined as \[n!=1 \times 2 \times 3 \times 4 \times 5....\times n\] so \[2!=1\times 2=2\] \[3!=1\times 2 \times 3=6\] Do you understand this?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

:D good, tell me what's 5! ?

OpenStudy (ash2326):

@Lunalovegood are you trying?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

let me know if you get stuck, whenever you say we'll proceed with the problem

OpenStudy (anonymous):

i'm stuck

OpenStudy (ash2326):

Ok, did you use the definition of n! to find 5!?

OpenStudy (anonymous):

yes,

OpenStudy (ash2326):

Where have you got stuck?

OpenStudy (anonymous):

ummm...well you could say like the first part

OpenStudy (ash2326):

\[5!=1 \times 2 \times 3 \times 4 \times 5\] or \[4!=5\times 4 \times 3 \times 3 \times 1\] Do you get this?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

Now you know \[^A C_B =\frac{A!}{B!(A-B)!}\] Can yo find 2C2

OpenStudy (anonymous):

@ash2326 , I will take over .:D

OpenStudy (ash2326):

@Lunalovegood I gotta go now. Sorry:( My friend @shivam_bhalla will help you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok. Yes @Lunalovegood , tell me what you got for 2C2

OpenStudy (anonymous):

@Lunalovegood , are you there ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did you find 2C2 or are you facing any problems ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Lunalovegood , yes for what ??

OpenStudy (anonymous):

i' am facing problems

OpenStudy (anonymous):

See, \[2C_2 = { 2! \over (2-2)! 2!}\] Now @Lunalovegood , you should know that 0! = 1 Rest you can solve. Now tell me what is 2C2 = ??

OpenStudy (anonymous):

@Lunalovegood , still any problem ??

OpenStudy (anonymous):

@Lunalovegood , there ??